Show that line integral is independent of path by finding a potential function for F

F(x,y) = (sin x − x sin y) j + (cos y + y cos x) i


1
Expert's answer
2021-07-09T07:00:53-0400

Let "P(x, y)=\cos y+y\cos x" and "Q(x, y)=\sin x-x\sin y." Then


"\dfrac{\partial P}{\partial y}=-\sin y+\cos x"

"\dfrac{\partial Q}{\partial x}=\cos x-\sin y"

"\dfrac{\partial P}{\partial y}=\cos x-\sin y=\dfrac{\partial Q}{\partial x}"

Then "F(x, y)" is conservative.


"\exist f , \nabla f=\vec F"


"\dfrac{\partial f}{\partial x}=P(x, y)=\cos y+y\cos x"

"\dfrac{\partial f}{\partial y}=Q(x, y)=\sin x-x\sin y"

"f=\int(\cos y+y\cos x)dx+\varphi(y)"

"=x\cos y+y\sin x+\varphi(y)"

"\dfrac{\partial f}{\partial y}=-x\sin y+\sin x+\varphi'(y)=\sin x-x\sin y"

"=>\varphi'(y)=0=>\varphi(y)=C"

"f(x, y)=x\cos y+y\sin x+C"

Therefore the line integral is independent of path.